Circular No. 8201

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003hz, 2003ia, 2003ib, 2003ic, 2003id Further to IAUC 8199, D. Singer, B. Beutler, B. Swift, and W. Li

report the LOTOSS discovery of four apparent supernovae (2003hz, 2003ib, 2003ic, 2003id) on unfiltered KAIT images. Also, H. Yamaoka, Kyushu University, reports the discovery by K. Itagaki, Yamagata, Japan, of an apparent supernova (SN 2003ia in NGC 6109) on unfiltered CCD images taken on Sept. 15.422 UT with a 0.60-m reflector.

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003hz Sept. 14.5  5^{h}48^{m}07.^{s}86 +46^{o}15′22′′.8 16.7 2′′.6 W, 1′′.6 N
2003ia Sept. 15.422 16 17 41.17 +35 00 07.2 17.2 7′′.5 E, 8′′.2 S
2003ib Sept. 16.2 20 33 30.98 −24 37 15.0 17.6 1′′.0 E, 10′′.4 N
2003ic Sept. 16.4  0 41 50.23 − 9 18 19.0 17.6 2′′.1 W, 7′′.6 S
2003id Sept. 16.5  2 21 35.70 − 5 31 51.0 16.7 11′′.3 W, 34′′.6 S

Additional approximate KAIT magnitudes: SN 2003hz in PGC 17866, Mar. 11.2, [19.5; Sept. 15.5, 16.8. SN 2003ib in MCG –04-48-15, Aug. 15.3, [18.5; Sept. 1.2, 19.0; 11.2, 17.9; 17.2, 17.5. SN 2003ic in MCG –02-2-86, Sept. 2.4, [19.0; 17.4, 17.6. SN 2003id in NGC 895, Sept. 7.5, [18.5; 17.5, 16.7. Additional magnitudes by Itagaki for SN 2003ia: 2002 June 2, [18.5; 5, [18.0; Sept. 16.411, 17.3.

SUPERNOVA 2003hx IN NGC 2076 L. Wang, Lawrence Berkeley National Laboratory; and D. Baade,

European Southern Observatory (ESO), communicate: “We observed SN 2003hx (cf. IAUC 8199, 8200) with the ESO Very Large Telescope (+ FORS1) in spectropolarimetry mode on Sept. 15.4 UT. The spectroscopy shows SN 2003hx to be a type-Ia supernova, ∼ 10 days past maximum. The interstellar Na I D line has an equivalent width of 4.86, indicating significant dust extinction. The observed degree of polarization is ∼ 2 percent and peaks blueward of 400 nm with a postion angle of 3^{o}.4. If this polarization is due to dust in the host galaxy, it implies that the dust particles are significantly smaller in size than their Galactic counterparts. The ratio of total-to-selective extinction, R_{v} = A_{v}/E(B − V), inferred from these observations is around 2.2.”

2003 September 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green